From nch@roe.ac.uk Fri Nov 7 23:41:30 2008 Date: Fri, 7 Nov 2008 16:26:33 +0000 (GMT) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Brian Walshe , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 7/11/08 Minutes of WFAU VDFS Science Archive meeting: 7th November 2008 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, LGR, MAR, ETWS, RPB Apologies: JPE, JMS, AL, BCW, RGM, NJC, RSC, MSH, PMW DoNM: 10am, Friday 14th November 2008 in the Vista Hut Actions discharged this week: ----------------------------- ACTION: LGR/MAR to respond and take appropriate action concerning recent GPS support requests. Discharged (20GB of binaries now with the GPS folks for their cluster search) - thanks to LGR. ACTION: NCH to allocate 07B H2 frames to the GPS programme and then let MAR & ETWS know when it's done. Discharged; ETWS also ingested the associated catalogues into the GPS detection tables. ACTION: NCH to ask RSC to check public/load-server synchronisation for robustness against load-server downtime. Discharged; RSC noted: "Yes, that was the first thing I implemented prior to letting it run live. Since then it has demonstrably worked on many occasion in the past when the WSA was down. The only time it might scrub the old database prior to completing the new one is if something goes wrong between recreating the new database and finishing table loads. Some of these things I've spotted and put checks in prior to recreation of the database. Although its impossible to cover every possible eventually up front, the case of the WSA being down was an easy one to solve." Actions partly discharged but continuing: ----------------------------------------- ACTION: NCH to remember to ask ETWS to update the GLIMPSE tables for GPS cross-matching in DR5. Nearly done, thanks to ETWS. Just need to reinitiate the WSA curation to get the updates into the WSA - NCH will show RPB how to do this. Actions carried forward from 31/10/08 meeting: ---------------------------------------------- ACTION: NCH to break everything by deprecating the loadServer DbConstant so that curations codes can be generalised to work for both VSA and WSA... but not until the WSA non-survey SW has been rolled out. - continues ACTION: All to comment on the development of the VSA metadata schema as described on the relevant TWiki page - continues ACTION: ETWS to add a delta-mag attribute to the VSA template detection table schema. - continues Specific points and new actions: -------------------------------- Project management: Nothing of note this week. WFCAM & VISTA updates: Nothing to report this week. Comments and issues arising from CASU fortnightly minutes: No new minutes this week. Networking: NJC and MJI have been chatting over beer and moose burgers in Quebec. Some discussions concerning extended source photometric attributes, and now seems like a good opportunity to tweak the standard catalogue output for VISTA. Suggestion is to substitute half-light flux and radius for the (seemingly unused) Hall values, and also a move to elliptical apertures. NCH noted we must make corresponding changes to the template VSA catalogue schema to keep in line. Concerning the extended, deblended source corrections, NCH expressed the concern that routine transfer/ingest ops for 08B could be held up waiting for the fix. ACTION: NCH to contact MJI to suggest continuation of routine ops for 08B despite the known funnies in the deblends. WSA Operations: RPB reported that load server ahmose has been misbehaving this week, and needed a replacement RAID card. This stopped the planned JBOD upgrade from going ahead (see below). ETWS noted: - Updated the sql scripts of the relevant tables for GLIMPSEI and GLIMPSEII. - Enhanced CU4 for ingestion of data into a different than the observed survey and processed and ingested U/07B/D3 into the GPS. Hardware MSH reported: "The network issues we experienced were due to nfs mount problems as all the linux servers were trying to mount ahmose's share drive, but this has been fixed as it has been unmounted from all (except khafre as someone seems to have been logged in and cd'd into the share preventing it from unmounting). The first 7 nodes of the cluster have been delivered, and the final node will be delivered once Eclipse finish testing the Windows 2008 drivers for the infiniband cards." RPB noted that the proposed disk upgrade on ahmose needs a little more thought since it turns out that the JBODs in question are still under maintenance, while 2 on the public server are no longer. NCH suggested that we should upgrade where we need the space most urgently and not worry too much about residual maintenance contracts, but in any case RPB will think on it and we'll take a decision next week with PMW's advice. ITSG have informed us that things are progressing with new server room C2: A/C is up and running and network infrastructure all in. Just need some special power connectors apparently. Software: ETWS noted: - Installed Python 2.6 on khafre and horus (64bit) and it seems to work better than the 2.5 release. - Also installed the latest version of HEASOFT and CFITSIO on khafre and horus and imcopy is working fine now. MAR noted that the colour image facilities in the interface are nearly ready to roll out. Survey Data Release: MAR noted that DR5 QC will start next week. Non-survey Data Release: Nothing to report this week Astrogrid deployment & Data Analysis services Concerning the new Intra-UKIDSS magnitude tool, LGR noted work on retrieval of filenames by rearranging one of MAR's web tools; applying illumination correction; investigating discrepancy with catalogued data (possibly due to parameters involved in the flux measurement). Miscellaneous: Nothing else to report this week. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================